The answer to ∬D ysin(xy)dA where D={(x,y)∈R^2:1≤x≤2,0≤y≤π}

is
Select one:
a. 0
b. (3π)/2


c. π/2

d. π

e. 1
1
Expert's answer
2020-06-04T21:19:17-0400

"\begin{aligned}
&\iint y \sin (xy)dA\textcolor{blue}{=\int_{0}^{\pi}\textcolor{red}{ \int_{1}^{2} y \sin (xy) dx}dy }\\[1 em]

\textcolor{red}{ \int_{1}^{2} y \sin (xy)dx}& =\frac{-y}{y}\cos (xy)\mid _{1}^{2} =-\cos (xy)\mid _{1}^{2}\\[1 em]
 &=-(\cos 2y -\cos y) =-\cos 2y +\cos y\\[1 em]
\textcolor{blue}{ \int_{0}^{\pi} -\cos 2y +\cos y dy} &=\frac{-1}{2}\sin 2y+\sin y\mid _{0}^{\pi}\\[1 em]
 &=(\frac{-1}{2}\sin 2\pi+\sin \pi)-(\frac{-1}{2}\sin 0+\sin 0)=\fcolorbox{red}{aqua}{0}\\[1 em]
\text{The answer is }\fcolorbox{red}{aqua}{a}
\end{aligned}"